OpenStudy (anonymous):

I have a simple question #17 \[\left| 4-3x \right|=3x-4\] what is the set of answers of x?

sam (.sam.):

|4-3x|=3x-4 4-3x=+-(3x-4) 4-3x=3x-4 4-6x=-4 -6x=-8 x=4/3

sam (.sam.):

For the 4-3x=-(3x-4) I don't think you get any value for x

OpenStudy (dumbcow):

4-3x = -(3x-4) 4-3x = -3x +4 0=0 Any x is a solution, so set is all real numbers

OpenStudy (anonymous):

cant be all real number try x=1

OpenStudy (accessdenied):

if you graph it, you can see that the right-hand side of the absolute value graph is the same line as the line it is equal to. the left-hand side goes off in another direction.

OpenStudy (dumbcow):

oh yeah, only if 3x-4 is positive x>= 4/3